# Stein spaces and Stein algebras

### v1

@@ Introduction
(0.1) { Stein spacesand holomorphic maps}{ Stein algebrasand -algebra morphisms} $\left\{\begin{array}[]{l}\textrm{\hskip 22.0ptStein spaces}\\ \textrm{and holomorphic maps}\end{array}\right\}\to\left\{\begin{array}[]{l}\textrm{\hskip 25.00003ptStein algebras}\\ \textrm{and $\mathbb{C}$-algebra morphisms}\end{array}\right\}$
@@ 1. Tools from Oka theory
$Y_{r}:=\{(z_{1},z_{2})\in\mathbb{C}^{2}\mid\mathrm{Im}(z_{2})<|z_{1}|^{2}+\mathrm{Re}(z_{2})^{2}+r\}.$
@@
Fix $r\in\mathbb{R}$. There is a homotopy $(h_{t})_{t\in[0,1]}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ inducing strong deformation retractions of both $\mathbb{C}^{2}$ and $Y_{r}$ onto $\{(z_{1},z_{2})\in\mathbb{C}^{2}\mid\mathrm{Im}(z_{2})\leq r-1\}$.
@@
[…]
@@
Define $U:=f_{1}^{-1}(Y_{r})$. It is an open neighborhood of $S^{\prime}$ in $S$. Let $Z\subset U$ be a closed neighborhood of $S^{\prime}$ in $U$. By the Tietze–Urysohn extension theorem, there exists a continuous map $\tau:S\to[0,1]$ which is equal to 00 on $Z$ and to $1$ on $S\setminus U$.
@@ 2. Holomorphic maps with finite-dimensional fibers
Fix $s\in S$. Write $A=\mathcal{O}_{S,s}$ and let $I_{1}$ (resp.$I_{2}$) be the ideal of $A$ consisting of germs of functions vanishing on $S_{1}$ (resp.$S_{2}$). Then the exactness of (2.1) at $s$ results from the exactness of $A\to A/I_{1}\oplus A/I_{2}\to A/\langle I_{1},I_{2}\rangle\to 0$, which is valid for any two ideals $I_{1}$ and $I_{2}$ of a commutative ring $A$.
@@ 3. Morphisms of Stein algebras
(3.1) { Stein spacesand holomorphic maps}{ Stein algebrasand -algebra morphisms} $\left\{\begin{array}[]{l}\textrm{\hskip 22.0ptStein spaces}\\ \textrm{and holomorphic maps}\end{array}\right\}\to\left\{\begin{array}[]{l}\textrm{\hskip 25.00003ptStein algebras}\\ \textrm{and $\mathbb{C}$-algebra morphisms}\end{array}\right\}$
@@
We finally record the following consequence of Theorem 3.3 for later use in [Ben]. If $S$ is a Stein space, we let $\lambda_{S}:S\to\mathrm{Spec}(\mathcal{O}(S))$ be the unique morphism of locally ringed spaces such that $\lambda_{S}^{*}:\mathcal{O}(S)\to\mathcal{O}(S)$ is the identity (see [SP, Lemma 01I1]).
@@
(3.2) {holomorphic maps XS}{morphisms of -locally ringed spaces $\left\{\begin{array}[]{l}\textrm{holomorphic maps}\\ \textrm{\hskip 23.99997pt$X\to S$}\end{array}\right\}\to\left\{\begin{array}[]{l}\textrm{morphisms of $\mathbb{C}$-locally ringed spaces}\\ \textrm{\hskip 38.99998pt$X\to\mathrm{Spec}(\mathcal{O}(S))$}\end{array}\right\}$

### v2

@@ Introduction
(0.1) { $\left\{\begin{array}[]{l}\textrm{\hskip 22.48872ptStein spaces}\\ \textrm{and holomorphic maps}\end{array}\right\}\to\left\{\begin{array}[]{l}\textrm{\hskip 25.55542ptStein algebras}\\ \textrm{and $\mathbb{C}$-algebra morphisms}\end{array}\right\}$
@@ 1. Tools from Oka theory
$Y_{r}:=\{(z_{1},z_{2})\in\mathbb{C}^{2}\mid\mathrm{Im}(z_{2})<|z_{1}|^{2}+\mathrm{Re}(z_{2})^{2}-r\}.$
@@
Fix $r\in\mathbb{R}$. There is a homotopy $(h_{t})_{t\in[0,1]}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ inducing strong deformation retractions of both $\mathbb{C}^{2}$ and $Y_{r}$ onto $\{(z_{1},z_{2})\in\mathbb{C}^{2}\mid\mathrm{Im}(z_{2})\leq-r-1\}$.
@@
$h_{t}(z_{1},z_{2})=(z_{1},z_{2}-it(\mathrm{Im}(z_{2})+r+1))$  $\textrm{ \hskip 10.00002ptif }\mathrm{Im}(z_{2})\geq-r-1$
$h_{t}(z_{1},z_{2})=(z_{1},z_{2})$  $\textrm{ \hskip 10.00002ptif }\mathrm{Im}(z_{2})\leq-r-1$
@@
Define $U:=f_{1}^{-1}(Y_{r})$. It is an open neighborhood of $S^{\prime}$ in $S$. Let $Z\subset U$ be a closed neighborhood of $S^{\prime}$ in $U$. By the Tietze–Urysohn extension theorem, there exists a continuous map $\tau:S\to[0,1]$ which is equal to 0 on $Z$ and to $1$ on $S\setminus U$.
@@ 2. Holomorphic maps with finite-dimensional fibers
[…]
@@ 3. Morphisms of Stein algebras
(3.1) { $\left\{\begin{array}[]{l}\textrm{\hskip 22.48872ptStein spaces}\\ \textrm{and holomorphic maps}\end{array}\right\}\to\left\{\begin{array}[]{l}\textrm{\hskip 25.55542ptStein algebras}\\ \textrm{and $\mathbb{C}$-algebra morphisms}\end{array}\right\}$
@@
We finally record the following consequence of Theorem 3.3 for use in [Ben24]. If $S$ is a Stein space, we let $\lambda_{S}:S\to\mathrm{Spec}(\mathcal{O}(S))$ be the unique morphism of locally ringed spaces such that $\lambda_{S}^{*}:\mathcal{O}(S)\to\mathcal{O}(S)$ is the identity (see [SP, Lemma 01I1]).
@@
(3.2) $\left\{\begin{array}[]{l}\textrm{holomorphic maps}\\ \textrm{\hskip 24.53313pt$X\to S$}\end{array}\right\}\to\left\{\begin{array}[]{l}\textrm{morphisms of $\mathbb{C}$-locally ringed spaces}\\ \textrm{\hskip 39.86638pt$X\to\mathrm{Spec}(\mathcal{O}(S))$}\end{array}\right\}$